OpenStudy (anonymous):

need help solve the system of equations by graphing then classify the system as consistent or inconsistent, and as dependent or independent 4x-y=15 x+5y=-12

OpenStudy (anonymous):

inconsistent x=21/2 y=-9/2

OpenStudy (anonymous):

would you plz show your work so i can figure out the next one on my own

OpenStudy (anonymous):

by graphing, you can find the point of intersection as the solution.

OpenStudy (anonymous):

will you plz show me how you got this

OpenStudy (anonymous):

first you plot the graph of these linear equations. they are straight lines. 4x-y=15 is y=4x-15 which 15 is the y-intercept. to find the x-intercept, substitue y=0 into the equation. 0=4x-15 implies x=15/4. now u can draw the straight line. same goes on x+5y=-12. change it to y=mx+c then you'll get the y-intercept. to find x-intercept, substitute y=0. after that plot them and you'll get another straight line. after plotting, you can see there is a point of intersection which is the solution of this linear system.

OpenStudy (anonymous):

i dont understand what the answer is what numbers do i plot

OpenStudy (anonymous):

plot the x-intercept and y-intercept that you obtained and draw a line pass through the points :)

OpenStudy (anonymous):

what are they? will you show your work

OpenStudy (anonymous):

for 4x-y=15, you have x-intercept = 15/4 and y-intercept = 15. put a dot on x=15/4 on x-axis and put a dot on y=15 at y-axis.

OpenStudy (anonymous):

then draw a straight line that pass through both of the dots. okay?

OpenStudy (anonymous):

not sure how to plot 15/4

OpenStudy (anonymous):

15/4 = 3 and 3/4. plot it in between 3 and 4 but close to 4 since the ratio is 3/4.

OpenStudy (anonymous):

so it would be (0,3/4)?

OpenStudy (anonymous):

15/4 is the x-intercept, you should draw it on x-axis. the coordinate is (15/4, 0). 15/4 is an improper fraction. you can change it into mixed number and you'll get 3+3/4 which is between 3 and 4 :)

OpenStudy (anonymous):

ok so i put a dot on (0,15) and between 3 and 4 what about the 2nd equation of x+5y=-12

OpenStudy (anonymous):

you're right. now we proceed to the next line. try to rearrange it into general form, y=mx+c.

OpenStudy (anonymous):

5y=mx+-12?

OpenStudy (anonymous):

5y=-x-12, divide all by 5, you'll get y=-x/5 -12/5. now you know the y-intercept is -12/5. try substituting y=0 and find the x value. that's how you obtain the x-intercept.